\begin{thebibliography}{10}\itemsep=-1pt

\bibitem{bai2021transformers}
Yutong Bai, Jieru Mei, Alan~L Yuille, and Cihang Xie.
\newblock Are transformers more robust than cnns?
\newblock {\em Advances in Neural Information Processing Systems},
  34:26831--26843, 2021.

\bibitem{bevandic2020multi}
Petra Bevandi{\'c}, Marin Or{\v{s}}i{\'c}, Ivan Grubi{\v{s}}i{\'c}, Josip
  {\v{S}}ari{\'c}, and Sini{\v{s}}a {\v{S}}egvi{\'c}.
\newblock Multi-domain semantic segmentation with pyramidal fusion.
\newblock {\em arXiv preprint arXiv:2009.01636}, 2020.

\bibitem{caesar2018coco}
Holger Caesar, Jasper Uijlings, and Vittorio Ferrari.
\newblock Coco-stuff: Thing and stuff classes in context.
\newblock In {\em Proceedings of the IEEE conference on computer vision and
  pattern recognition}, pages 1209--1218, 2018.

\bibitem{chen2017deeplab}
Liang-Chieh Chen, George Papandreou, Iasonas Kokkinos, Kevin Murphy, and Alan~L
  Yuille.
\newblock Deeplab: Semantic image segmentation with deep convolutional nets,
  atrous convolution, and fully connected crfs.
\newblock {\em IEEE transactions on pattern analysis and machine intelligence},
  40(4):834--848, 2017.

\bibitem{contributors2020mmsegmentation}
MMSegmentation Contributors.
\newblock Mmsegmentation: Openmmlab semantic segmentation toolbox and
  benchmark.
\newblock {\em Availabe online: https://github. com/open-mmlab/mmsegmentation
  (accessed on 18 May 2022)}, 2020.

\bibitem{cordts2016cityscapes}
Marius Cordts, Mohamed Omran, Sebastian Ramos, Timo Rehfeld, Markus Enzweiler,
  Rodrigo Benenson, Uwe Franke, Stefan Roth, and Bernt Schiele.
\newblock The cityscapes dataset for semantic urban scene understanding.
\newblock In {\em Proceedings of the IEEE conference on computer vision and
  pattern recognition}, pages 3213--3223, 2016.

\bibitem{dai2017scannet}
Angela Dai, Angel~X Chang, Manolis Savva, Maciej Halber, Thomas Funkhouser, and
  Matthias Nie{\ss}ner.
\newblock Scannet: Richly-annotated 3d reconstructions of indoor scenes.
\newblock In {\em Proceedings of the IEEE conference on computer vision and
  pattern recognition}, pages 5828--5839, 2017.

\bibitem{goyal2017accurate}
Priya Goyal, Piotr Doll{\'a}r, Ross Girshick, Pieter Noordhuis, Lukasz
  Wesolowski, Aapo Kyrola, Andrew Tulloch, Yangqing Jia, and Kaiming He.
\newblock Accurate, large minibatch sgd: Training imagenet in 1 hour.
\newblock {\em arXiv preprint arXiv:1706.02677}, 2017.

\bibitem{hendrycks2018benchmarking}
Dan Hendrycks and Thomas Dietterich.
\newblock Benchmarking neural network robustness to common corruptions and
  perturbations.
\newblock In {\em International Conference on Learning Representations}, 2018.

\bibitem{lambert2020mseg}
John Lambert, Zhuang Liu, Ozan Sener, James Hays, and Vladlen Koltun.
\newblock Mseg: A composite dataset for multi-domain semantic segmentation.
\newblock In {\em Proceedings of the IEEE/CVF conference on computer vision and
  pattern recognition}, pages 2879--2888, 2020.

\bibitem{lin2014microsoft}
Tsung-Yi Lin, Michael Maire, Serge Belongie, James Hays, Pietro Perona, Deva
  Ramanan, Piotr Doll{\'a}r, and C~Lawrence Zitnick.
\newblock Microsoft coco: Common objects in context.
\newblock In {\em European conference on computer vision}, pages 740--755.
  Springer, 2014.

\bibitem{loshchilov2018decoupled}
Ilya Loshchilov and Frank Hutter.
\newblock Decoupled weight decay regularization.
\newblock In {\em International Conference on Learning Representations}, 2018.

\bibitem{naseer2021intriguing}
Muhammad~Muzammal Naseer, Kanchana Ranasinghe, Salman~H Khan, Munawar Hayat,
  Fahad Shahbaz~Khan, and Ming-Hsuan Yang.
\newblock Intriguing properties of vision transformers.
\newblock {\em Advances in Neural Information Processing Systems},
  34:23296--23308, 2021.

\bibitem{neuhold2017mapillary}
Gerhard Neuhold, Tobias Ollmann, Samuel Rota~Bulo, and Peter Kontschieder.
\newblock The mapillary vistas dataset for semantic understanding of street
  scenes.
\newblock In {\em Proceedings of the IEEE international conference on computer
  vision}, pages 4990--4999, 2017.

\bibitem{paul2022vision}
Sayak Paul and Pin-Yu Chen.
\newblock Vision transformers are robust learners.
\newblock In {\em Proceedings of the AAAI Conference on Artificial
  Intelligence}, volume~36, pages 2071--2081, 2022.

\bibitem{richter2017playing}
Stephan~R Richter, Zeeshan Hayder, and Vladlen Koltun.
\newblock Playing for benchmarks.
\newblock In {\em Proceedings of the IEEE International Conference on Computer
  Vision}, pages 2213--2222, 2017.

\bibitem{varma2019idd}
Girish Varma, Anbumani Subramanian, Anoop Namboodiri, Manmohan Chandraker, and
  CV Jawahar.
\newblock Idd: A dataset for exploring problems of autonomous navigation in
  unconstrained environments.
\newblock In {\em 2019 IEEE Winter Conference on Applications of Computer
  Vision (WACV)}, pages 1743--1751. IEEE, 2019.

\bibitem{wang2022generalizing}
Jindong Wang, Cuiling Lan, Chang Liu, Yidong Ouyang, Tao Qin, Wang Lu, Yiqiang
  Chen, Wenjun Zeng, and Philip Yu.
\newblock Generalizing to unseen domains: A survey on domain generalization.
\newblock {\em IEEE Transactions on Knowledge and Data Engineering}, 2022.

\bibitem{xie2021segformer}
Enze Xie, Wenhai Wang, Zhiding Yu, Anima Anandkumar, Jose~M Alvarez, and Ping
  Luo.
\newblock Segformer: Simple and efficient design for semantic segmentation with
  transformers.
\newblock {\em Advances in Neural Information Processing Systems},
  34:12077--12090, 2021.

\bibitem{yu2020bdd100k}
Fisher Yu, Haofeng Chen, Xin Wang, Wenqi Xian, Yingying Chen, Fangchen Liu,
  Vashisht Madhavan, and Trevor Darrell.
\newblock Bdd100k: A diverse driving dataset for heterogeneous multitask
  learning.
\newblock In {\em Proceedings of the IEEE/CVF conference on computer vision and
  pattern recognition}, pages 2636--2645, 2020.

\bibitem{zendel2018wilddash}
Oliver Zendel, Katrin Honauer, Markus Murschitz, Daniel Steininger, and
  Gustavo~Fernandez Dominguez.
\newblock Wilddash-creating hazard-aware benchmarks.
\newblock In {\em Proceedings of the European Conference on Computer Vision
  (ECCV)}, pages 402--416, 2018.

\bibitem{zhou2017scene}
Bolei Zhou, Hang Zhao, Xavier Puig, Sanja Fidler, Adela Barriuso, and Antonio
  Torralba.
\newblock Scene parsing through ade20k dataset.
\newblock In {\em Proceedings of the IEEE conference on computer vision and
  pattern recognition}, pages 633--641, 2017.

\bibitem{zhou2022understanding}
Daquan Zhou, Zhiding Yu, Enze Xie, Chaowei Xiao, Animashree Anandkumar, Jiashi
  Feng, and Jose~M Alvarez.
\newblock Understanding the robustness in vision transformers.
\newblock In {\em International Conference on Machine Learning}, pages
  27378--27394. PMLR, 2022.

\bibitem{zhou2022domain}
Kaiyang Zhou, Ziwei Liu, Yu Qiao, Tao Xiang, and Chen~Change Loy.
\newblock Domain generalization: A survey.
\newblock {\em IEEE Transactions on Pattern Analysis and Machine Intelligence},
  2022.

\end{thebibliography}

\documentclass[10pt,twocolumn,letterpaper]{article}
\pdfoutput=1

\usepackage[pagenumbers]{cvpr} 

\usepackage{graphicx}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{booktabs}

\usepackage{graphicx, amsmath, amssymb, caption, multirow, overpic, textpos}
\usepackage[table]{xcolor}
\usepackage[british, american]{babel}
\definecolor{citecolor}{RGB}{119,185,0}
\definecolor{linkcolor}{HTML}{ED1C24}
\usepackage[pagebackref=false, breaklinks=true, colorlinks,
            citecolor=citecolor, linkcolor=linkcolor, bookmarks=false]{hyperref}

\newlength\savewidth\newcommand\shline{\noalign{\global\savewidth\arrayrulewidth
  \global\arrayrulewidth 1pt}\hline\noalign{\global\arrayrulewidth\savewidth}}
\newcommand{\tablestyle}[2]{\setlength{\tabcolsep}{#1}\renewcommand{\arraystretch}{#2}\centering\footnotesize}
\renewcommand{\paragraph}[1]{\vspace{1.25mm}\noindent\textbf{#1}}

\newcolumntype{x}[1]{>{\centering\arraybackslash}p{#1pt}}
\newcolumntype{y}[1]{>{\raggedright\arraybackslash}p{#1pt}}
\newcolumntype{z}[1]{>{\raggedleft\arraybackslash}p{#1pt}}

\usepackage[capitalize]{cleveref}
\crefname{section}{Sec.}{Secs.}
\Crefname{section}{Section}{Sections}
\Crefname{table}{Table}{Tables}
\crefname{table}{Tab.}{Tabs.}

\begin{document}

\title{1st Place Solution of The Robust Vision Challenge 2022 \\
Semantic Segmentation Track}

\author
{
Junfei Xiao${^{1}}$
\quad
Zhichao Xu${^{3}}$
\quad
Shiyi Lan${^{2}}$
\quad
Zhiding Yu${^{2}}$
\quad
Alan Yuille${^{1}}$
\quad
Anima Anandkumar${^{2}}$
\\$^{1}$Johns Hopkins University \qquad ~$^{2}$NVIDIA \qquad ~$^{3}$Fudan University
}
\maketitle

\begin{abstract}
   This report describes the winning solution to the Robust Vision Challenge (RVC) semantic segmentation track at ECCV 2022. Our method adopts the FAN-B-Hybrid model as the encoder and uses SegFormer as the segmentation framework. The model is trained on a composite dataset consisting of images from 9 datasets (ADE20K, Cityscapes, Mapillary Vistas, ScanNet, VIPER, WildDash 2, IDD, BDD, and COCO) with a simple dataset balancing strategy. All the original labels are projected to a 256-class unified label space, and the model is trained using a cross-entropy loss. Without significant hyperparameter tuning or any specific loss weighting, our solution ranks the first place on all the testing semantic segmentation benchmarks from multiple domains (ADE20K, Cityscapes, Mapillary Vistas, ScanNet, VIPER, and WildDash 2). The proposed method can serve as a strong baseline for the multi-domain segmentation task and benefit future works. Code will be available at \url{https://github.com/lambert-x/RVC_Segmentation}.
\end{abstract}

\section{Introduction}
\label{sec:intro}

In the past few years, advances in deep learning have led to significant progress in visual recognition. However, the robustness of state-of-the-art deep learning models remains an open issue. On the one hand, real-world applications require models to be deployed ``in the wild''. On the other hand, many current deep models have been shown to be brittle to distributional shifts and natural perturbations. This phenomenon raised considerable interest in open problems such as domain generalization and adaptation.

There is rich literature in domain generalization~\cite{wang2022generalizing,zhou2022domain} where popular methods include, but are not limited to: domain randomization, domain invariant representation learning, disentanglement learning and meta learning, etc. One approach related to this work is multi-dataset training~\cite{lambert2020mseg}, in which the authors show that a simple combination of multiple datasets with label space alignment can outperform strong domain generalization approaches.

Another interesting trend is the recent surge of Vision Transformers (ViTs). Several works~\cite{paul2022vision,naseer2021intriguing,bai2021transformers,xie2021segformer} almost simultaneously pointed out that ViTs demonstrate surprisingly strong robustness to out-of-distribution scenarios. For example, SegFormer~\cite{xie2021segformer} demonstrates significantly better results over CNN-based strong methods in Cityscapes-C, a more challenging variant of Cityscapes contaminated by 16 types of natural corruption. More recently, \cite{zhou2022understanding} introduced the fully attentional network (FAN), a family of ViT backbones with state-of-the-art accuracy and robustness in both image classification and downstream tasks.

This report describes the winning solution to the RVC 2022 semantic segmentation track. This year, the challenge features benchmarking of a single semantic segmentation model on six datasets, spanning both indoor/outdoor and synthetic/real. Thus, it presents a great challenge to the generalization capability of a model over different domains. Our solution is inspired by the above advances in both multi-dataset training and ViTs, as will be detailed in the rest of the report.

\section{Method}
\label{sec:method}

\begin{table}[t]
\centering
\small
\begin{tabular}{llcc}
\multirow{2}{*}{Dataset} & \multirow{2}{*}{Scenes} 
  & {\#Images }
  & {\#Class}

  \\
  & &  (train/val) & (orig-proj) \\
\hline
COCO~\cite{caesar2018coco,lin2014microsoft}       & Natural   &  118287/5000
  & 201$\rightarrow$133\\
ADE20K~\cite{zhou2017scene}       & Natural   &  20210/2000
  & 151$\rightarrow$146 \\
Cityscapes~\cite{cordts2016cityscapes}   & Driving  &   2975/500 
  & 34$\rightarrow$31    \\
Vistas~\cite{neuhold2017mapillary}       & Driving  &  18000/2000 
  &  66$\rightarrow$64 \\
BDD~\cite{yu2020bdd100k}   & Driving  &   7000/1000 
  & 19$\rightarrow$19     \\
  
IDD~\cite{varma2019idd}   & Driving  &   6993/981 
  & 39$\rightarrow$26    \\
WildDash 2~\cite{zendel2018wilddash}     & Driving  &   3413/857 
  & 34$\rightarrow$31     \\

ScanNet~\cite{dai2017scannet}      & Indoor &  19466/5436 
  & 41$\rightarrow$41    \\
VIPER~\cite{richter2017playing}        & Artificial &  13367/4959 
  & 32$\rightarrow$32    \\

\end{tabular}
\caption{\textbf{Datasets overview.}
  A total of 9 datasets across natural, driving, indoor, and artificial scenes are used for training and validating the model.
  Class count denotes the number of classes in the original label space
  and the projected label space.
  }
\label{tab:datasets}
\end{table}

\begin{table*}[!t]
\centering

 \tablestyle{8pt}{1.05}
\begin{tabular}{@{}lcccccccc@{}}
\toprule
  & \multicolumn{6}{c}{ \textsc{RVC Test Datasets} } \\
\textsc{Method Name} &\textsc{Year/Rank}   &  \textsc{ADE20K} & \textsc{Cityscapes}   & \textsc{Mapillary} & \textsc{ScanNet} & \textsc{VIPER} & \textsc{WildDash 2}  \\
\midrule

\textsc{MSeg1080\_RVC} \cite{lambert2020mseg} & {2020 / 2nd} & {{33.18}}   & {{80.7}}  & {34.19}   & {48.5}   & {40.7}  & {34.71} \\
\textsc{SN\_RN152pyrx8\_RVC} \cite{bevandic2020multi} & {2020 / 1st} & {31.12}   & {74.7}  & {{40.43}}     & {{54.6}}      & {{62.5}}  & {{42.29}}  \\

\textsc{FAN\_NV\_RVC} (Ours) & {2022 / 1st} & \textbf{43.46}   & \textbf{82.0}     & \textbf{{55.27}}     & \textbf{{58.6}}      & \textbf{{69.8}}  & \textbf{{47.5}}  \\

\bottomrule
\end{tabular}
 \caption{\textbf{Comparison with previous methods.}  Measured by class mIoU. The best number in each column is highlighted in bold.} 
\label{tab:results}
\end{table*}

\paragraph{Backbone.}
We adopt FAN-B-Hybrid~\cite{zhou2022understanding} as our backbone encoder due to its great robustness on multiple benckmarks (ImageNet-C~\cite{hendrycks2018benchmarking}, Cityscapes-C, etc.). The backbone is initialized with the weight pretrained on ImageNet-22K and fine-tuned on ImageNet-1K (the checkpoint is provided in the official github repository\footnote{\url{https://github.com/NVlabs/FAN}}).

\paragraph{Segmentation framework.}
We use SegFormer~\cite{xie2021segformer} as the segmentation framework. It uses simple but effective multilayer perceptron (MLP) decoders to fuse multi-level features (the outputs of the early Convolution blocks, last FAN Transformer block and the final class attention block output) and predict the semantic segmentation mask. The reader may refer to the official github of FAN (segmentation folder) for more details. Cross-entropy loss is used for training the model.

\paragraph{Training set.}
The model is trained on a composite dataset with all the training images from ADE20K, Cityscapes, Mapillary Vistas, ScanNet, VIPER, WildDash 2, IDD, BDD, and COCO. An overview of all the datasets involved is shown in \Cref{tab:datasets}. The datasets vary largely in size (COCO is more than 30 times larger than WildDash 2). To alleviate the imbalanced data size issue for better model generalization, we adopt a simple dataset resizing strategy - repeat each dataset $(120,000// len(dataset))$ times.

\paragraph{Unified label space.}
We directly use the unified label space provided in the official RVC github repository\footnote{\url{https://github.com/ozendelait/rvc_devkit}} (with some minor corrections), which has 256 classes. This label space is naive and noisy for relabeling some fine-grained classes in their original label space. 

\paragraph{Post-processing.} All predicted segmentation maps are projected from the unified label space to the original label space of every dataset.

\section{Implementation Details}
\label{sec:imple_details}
We built our codebase with MMSegmentation~\cite{contributors2020mmsegmentation}. The length of the training process is 80,000 iterations, while the first half training is without BDD and IDD datasets. \Cref{tab:imple_details} provides detailed information about the optimizer and hyperparameter settings. Training and testing data augmentations are detailed in \Cref{tab:train_aug} and \Cref{tab:test_aug}. The model is trained on 64 V100 GPUs (32G), and the whole training procedure takes $\sim$35 hours.

\begin{table}[ht]
\tablestyle{6pt}{1.02}
\small
\begin{tabular}{y{96}|y{68}}
Config & Setting \\
\shline
Optimizer & AdamW \cite{loshchilov2018decoupled} \\
Learning rate & 6e-5 \\
Weight decay & 0.01 \\
Optimizer momentum & $\beta_1, \beta_2{=}0.9, 0.999$ \\
Batch size & 64 \\
Learning rate schedule & Poly \cite{chen2017deeplab}\\
Warmup iters \cite{goyal2017accurate} & 1500 \\
\end{tabular}
\vspace{-.5em}
\caption{\textbf{Optimizer \& hyper-parameters details.}}
\label{tab:imple_details} \vspace{-.5em}
\end{table}

\begin{table}[ht]
\tablestyle{8pt}{1.05}
\small
\begin{tabular}{c|c}
 Operation & Setting \\ 
\shline
Resize &  Scale: (2048, 1024), Ratio: (0.5, 2.0) \\
RandomCrop &  Crop size: (1024, 1024) \\
RandomFlip & Prob: 0.5 \\
PhotoMetricDistortion & Default
\end{tabular}
\caption{\textbf{Training data augmentations.}}
\label{tab:train_aug} 
\vspace{-.5em}
\end{table}

\begin{table}[ht]
\tablestyle{12pt}{1.2}
\small
\begin{tabular}{c|c}
 Operation & Setting \\ 
\shline
 Resize & Scale: (2048, 1024) \\
 Multi-scale &  Ratios: (0.5, 0.75, 1.0, 1.25, 1.5, 1.75) \\
 Flip & True
\end{tabular}
\caption{\textbf{Testing data augmentations.}}
\label{tab:test_aug} 
\vspace{-.5em}
\end{table}

\section{Results}
\label{sec:results}

We compare our method with the winning solutions of RVC 2020 in \Cref{tab:results}. Our method makes solid improvements on all six benchmarks and beats them all by a large margin.

\section{Conclusion}
\label{sec:conclusion}

In this report, we described the winning solution of the RVC 2022 Semantic Segmentation Track. Our result shows that Vision Transformer models (in our case, FAN), when coupled with multi-dataset training, exhibit strong robustness and generalization at scales in semantic segmentation. Our work again echoes recent discoveries of improved robustness and representation in ViTs. However, it is worth noting that the training computation and memory consumption have become important challenges as the data and label space become large. The efficiency on devices during deployment also presents another challenge for real-world applications of current ViT models.

{\small
\bibliographystyle{ieee_fullname}
\bibliography{ref}
}

\end{document}